Convert the following real scores into standard scores.
Given: mean=9.25 and standard deviation= 2.14
a. x=10
b. x=12
c.x=6

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

Convert the following real scores into standard scores.
Given: mean = 9.25 and standard deviation= 2.14
---
Formula:
z(x) = (x-u)/sigma
-------
a. x=10
z(10) = (10-9.25)/2.14 = 0.75/2.14 = 0.35
-----------------------------------------------
b. x=12
z(12) = (12-9.25)/2.14 = 2.75/2.14 = 1.29
-----------------------------------------------
c.x=6
z(6) = (6-9.25)/2.14 = -1.52
===================================
Cheers,
Stan H.
============

Answer by solver91311(24713)





You can do your own arithmetic.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism